OpenStudy (anonymous):

Simplify this radical expression. Use absolute values when needed.... Answer choices: A) x^2+9 B) |x+3| C) x+3 D) |x^2+9|

OpenStudy (anonymous):

\[\sqrt{(x+3)^2}\]

jhonyy9 (jhonyy9):

so will be C). but can you tell me why ?

jhonyy9 (jhonyy9):

so can you writing please sqrt(x) = ? using exponent form

jhonyy9 (jhonyy9):

so you dont like cooperating anything please ?

jhonyy9 (jhonyy9):

so than sqrt(x)=x^(1/2) yes ?

OpenStudy (anonymous):

Im sorry i was trying to figure it out. i dont understand :/

jhonyy9 (jhonyy9):

so with this example i like to make you understanding all right hopw do you know that sqrt(x)=x^(1/2) where the denominator of fraction from the exponent always mean the indice of radical

jhonyy9 (jhonyy9):

so the same cuberoot(x)=x^(1/3) ok ?

jhonyy9 (jhonyy9):

so where is Zuni ? please

jhonyy9 (jhonyy9):

i am from Europe country Hungary

OpenStudy (anonymous):

okayy... i think i am kind of understanding what you are trying to say with the problem... Zuni is in the United States in the state of New Mexico :) Wow you are far far away!

jhonyy9 (jhonyy9):

so than your exercise will be using exponentialy form sqrt(x+3)^2 =((x+3)^2)^(1/2) so right ? do you understand it sure ?

jhonyy9 (jhonyy9):

and yes i am so far far away ,you see it right

OpenStudy (anonymous):

yes! :) !!

jhonyy9 (jhonyy9):

do you know it where is Hungary ?

OpenStudy (anonymous):

Is it in Eastern Europe?

jhonyy9 (jhonyy9):

yes aproximatly ,but is more precisly in midle of Europe

OpenStudy (anonymous):

Wow it must be very amazing living there!

jhonyy9 (jhonyy9):

you dont was again here never ? so but my birth country is Romania ,this is the east country of Hungary ,there is Transylvania what suppose you know if you like ,love the moves horror with Dracula

OpenStudy (anonymous):

I never seen the place. wow that is amazing! And i have never seen that movie but i like the movie Van Helsing

jhonyy9 (jhonyy9):

so ok but i like to help you solving your exercise ,ok ? than how you can rewriting it ?

OpenStudy (anonymous):

Okayy and what do you mean by rewriting it?

jhonyy9 (jhonyy9):

sorry was loss the internet conection so i have wrote it little up using the exponentialy form yes ? but there you can simplifie again

jhonyy9 (jhonyy9):

((x+3)^2)^(1/2) =(x+3)^((2*1)/2) =(x+3)^1 =x+3 ok ?

jhonyy9 (jhonyy9):

hope so much that you can understanding all right sure